OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1 = 1, an = 4an-1 - 1. Enter your answers separated by commas; example: 1, 5, 10, 15, 20. I got 3,11,43,171,683. Not sure what I'm doing wrong.

OpenStudy (anonymous):

Anyone? I've done the problem 3 times and following along with an example video and have no idea how this answer is wrong. We are using the recursive formula, so it should be 4(1)-1=3, then 4(3)-1=11, then 4(11)-1=43, and so on.

OpenStudy (anonymous):

But you said an=4an-1-1 so you need to subtract twice right?

OpenStudy (anonymous):

Ohh wait. You meant that's the first term.

OpenStudy (anonymous):

You forgot that the first term is 1 haha.

OpenStudy (anonymous):

So it's 1,3,11,43,171.

OpenStudy (anonymous):

Oh I tried that before but it said it was wrong because there weren't spaces between my answers. -_- This is the only time I've ever had to space my answers instead of just putting a comma, so I guess it was just a technical difficulty. Thanks though!

OpenStudy (anonymous):

You're welcome!